5x+5y=7 and 7x-3=13 how do i solve using elimination method?
Found 2 solutions by Alan3354, palanisamy:
Answer by Alan3354(69443)

5x+5y=7 and
7x-3=13
how do i solve using elimination method?
7x-3=13
7x = 16
x = 16/7
---------------
5(16/7) + 5y = 7
80/7 + 5y = 49/7
5y = -31/7
y = -31/35

Answer by palanisamy(496)

The given equations are,
5x+5y=7 . ..(1)
and 7x-3y=13 ...(2)
(1)*3 => 15x+15y = 21 ...(3)
(2)*5 => 35x-15y = 65 ...(4)
(3)+(4) => 50x = 86
x = 86/50
x = 43/25
(1) => 5*(43/25)+5y = 7
43/5+5y = 7
5y = 7-43/5
5y = (35-43)/5
y = -8/25

Therefore the solution is,
x = 43/25 and y = -8/25